OpenStudy (anonymous):

A.4/5 yd B. 9 yd C.5/5 yd D. 1/5 yd Angelo bought a chain that is 4 yards long. He wants to cut the chain into 5 equal-length pieces. How long will each piece of chain be?

OpenStudy (anonymous):

help!!!

OpenStudy (texaschic101):

4/5 yds or 0.8 yds

OpenStudy (anonymous):

can u help with more?

OpenStudy (texaschic101):

I can try...can't make no promises

OpenStudy (anonymous):

but the top one is right huh?

OpenStudy (texaschic101):

yes

OpenStudy (texaschic101):

there is 9 puppies...6 are black...that means 3 are brown....so the brown puppies make up 3/9 of the puppies

OpenStudy (anonymous):

so it would be D...

OpenStudy (texaschic101):

yes...it would be D

OpenStudy (anonymous):

can u help with more please

OpenStudy (texaschic101):

sure

OpenStudy (anonymous):

There are 4 members of the swim team practicing their backstroke. The other 6 team members are working on their butterfly stroke. What fraction of the swim team is practicing their backstroke? A. 4/10 B.4/6 C. 6/10 D. 10/4

OpenStudy (texaschic101):

practicing backstroke/ total swim team = 4/10

OpenStudy (anonymous):

here is another one ?

OpenStudy (texaschic101):

C

OpenStudy (anonymous):

r u sure

OpenStudy (texaschic101):

positive

OpenStudy (anonymous):

ok lets do another one ok

OpenStudy (texaschic101):

sure

OpenStudy (anonymous):

Tim is 7. His brother is 12. Which fraction represents Tim's age compared to his brother's age? A. 7/12 B. 5/12 C. 12/7 D. 5/7

OpenStudy (texaschic101):

pay attention to the order of the wording.. Tim to his brother would be 7/12 if it had said his brother to Tim, then it would have been 12/7

OpenStudy (anonymous):

i do pay attention i got 12/7

OpenStudy (texaschic101):

tims age compared to his brother would be 7/12

OpenStudy (anonymous):

so 7 /12 is the answer

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

i got 2 more can u help me ?

OpenStudy (texaschic101):

ask away

OpenStudy (texaschic101):

none of them do A = 2/8 or 1/4 B = 4/10 or 2/5 C = 4/6 or 2/3

OpenStudy (anonymous):

i has to be one of them

OpenStudy (texaschic101):

there must be a miss-print....because 2/4 is the same as 1/2....and no shape is half-shaded

OpenStudy (anonymous):

well here another

OpenStudy (texaschic101):

A

OpenStudy (texaschic101):

A shows 2 2/10 which is equal to 2 1/5

OpenStudy (anonymous):

so would it be A

OpenStudy (texaschic101):

yes...its A

OpenStudy (anonymous):

can i ask u another one ?

OpenStudy (texaschic101):

yes

OpenStudy (texaschic101):

2/5 and 8/20

OpenStudy (texaschic101):

3/5 * 5/5 = 15/25

OpenStudy (anonymous):

r u sure?

OpenStudy (texaschic101):

100% positive

OpenStudy (texaschic101):

3/8 and 4/10 ....common denominator is 40 15/40 and 16/40 4/10 is bigger.....he mowed more after lunch

OpenStudy (anonymous):

would it be C

OpenStudy (texaschic101):

yes

OpenStudy (texaschic101):

turn them to decimals... 7/8 = 0.875 3/4 = 0.75 8/12 = 0.666 9/10 = 0.9 10/15 = 0.666 answer is 9/10

OpenStudy (anonymous):

i just wanted to say thx for all of ur help but i 1 last one

OpenStudy (texaschic101):

your very welcome.....what is your last q ?

OpenStudy (texaschic101):

7/10 and 30/50 --- common denominator is 50 35/50 > 30/50 you would use > (greater then) 7/10 > 30/50

OpenStudy (anonymous):

final one i promise

OpenStudy (texaschic101):

D

OpenStudy (anonymous):

welll thx u if i need anymore help i'll take u alright i mean if that ok with u

OpenStudy (texaschic101):

no problem...I was glad to help. However, I have to go do something so I will be off for a little while. I will be back on later if you happen to need anymore help

OpenStudy (anonymous):

okay sound good see u later? and thx again

OpenStudy (texaschic101):

:)